# One-to-N Backdoor Attack in 3D Point Cloud via Spherical Trigger

### v1

@@ Abstract
Backdoor attacks represent a critical threat to deep learning systems, particularly in safety-sensitive 3D domains such as autonomous driving and robotics. However, existing backdoor attacks for 3D point clouds have been limited to a rigid one-to-one paradigm. To address this, we present the first one-to-N backdoor framework for 3D vision, based on a novel, configurable spherical trigger. Our key insight is to leverage the spatial properties of spheres as a parameter space, allowing a single trigger design to encode multiple target classes. We establish a theoretical foundation based on Neural Tangent Kernel theory for one-to-N backdoor attacks in 3D domain, demonstrating that poisoned models can map distinct trigger configurations to different target labels. Experimental results systematically validate this conclusion across multiple datasets and model architectures, achieving high attack success rates (up to 100%) while maintaining accuracy on clean data. This work establishes a crucial benchmark for multi-target threats in 3D vision and provides the foundational understanding needed to secure future 3D-driven intelligent systems.
@@ 1 Introduction
Current research on 3D backdoor attacks has primarily focused on one-to-one attack paradigms, where a single trigger pattern is designed to cause the model to misclassify any input containing it as a specific target class determined during the training phase Notable works in this domain include PointBA [4], which introduces orientation and interaction triggers for point clouds; IRBA [5], which employs weighted local transformations for imperceptible and robust attacks; and frequency-domain based approaches [6] that enhance stealthiness through spectral manipulations. While these methods have demonstrated impressive attack success rates, they share a fundamental limitation: the attack is confined at inference to misclassifying inputs into the sole target class determined during training. This constraint significantly reduces the attack flexibility and practicality in real-world scenarios where attackers may need to target multiple classes simultaneously.
@@
The concept of one-to-N backdoor attacks was first systematically formalized in the pioneering work by Xue et al. [7] for 2D image domains, representing a significant evolution beyond traditional one-to-one paradigms. This approach enables a single trigger design to map to $N$ target classes through parametric variations such as its intensity or spatial pattern. This contrasts sharply with the N-to-N paradigm of combining multiple distinct trigger types, which suffers from scalability limitations stemming from the finite number of available types, and performance inconsistencies due to divergent optimization requirements. Building upon this foundational concept, the one-to-N principle has been extended to more generalized settings [8, 9, 10, 11], including universal backdoor attacks [8] targeting all classes and Marksman backdoors [9] supporting arbitrary target selection during inference. However, the translation of this powerful one-to-N paradigm into the 3D domain presents distinct and unmet challenges. The structural transition from dense, grid-aligned pixels to unstructured point clouds invalidates core assumptions of 2D trigger design and demands a novel instantiation of the one-to-N principle, supported by theoretical analysis tailored to 3D geometric properties. Consequently, while existing 3D backdoor research has predominantly focused on one-to-one mappings, the potential for advanced one-to-N attacks has remained largely unexplored due to these inherent structural barriers.
To bridge this critical gap, we introduce the first dynamic one-to-N backdoor framework for 3D point clouds. Our key insight leverages simple spheres as a unified trigger entity, whose spatial properties (e.g., position, radius) form a flexible parameter space for encoding different target classes. A theoretical foundation based on Neural Tangent Kernel (NTK) theory is established to demonstrate the feasibility of such one-to-N mappings in 3D feature spaces, with experimental results systematically validating this theoretical conclusion across multiple datasets and model architectures. As the first systematic framework for one-to-N backdoor attacks in 3D, the proposed method achieves high attack success rates while offering significant flexibility.
The main contributions of this work are threefold:
@@
    First one-to-N backdoor framework for 3D point clouds: We pioneer a dynamic one-to-N backdoor attack framework for 3D point clouds, breaking the limitations of traditional one-to-one paradigms.
@@
The rest of this paper is organized as follows. Section 2 reviews related work on 3D point cloud deep learning, backdoor attacks in 2D images and 3D point clouds, and existing defense mechanisms. Section 3 introduces our proposed one-to-N backdoor attack framework, including the problem formulation, trigger design methodology, and attack mechanism. Section 4 presents comprehensive experimental results and analysis across multiple datasets and model architectures, evaluating attack effectiveness, scalability with varying target numbers, and resistance to defenses. Section 5 concludes the paper by summarizing key findings and discussing future research directions. Additionally, theoretical analysis supporting the feasibility of one-to-N mappings in 3D feature spaces and trigger placement algorithms are provided in the Appendix.
@@ 2.1 Deep Learning on 3D Point Cloud
The evolution of direct point cloud processing networks has addressed these challenges through several representative architectures. PointNet [12] pioneers the use of symmetric functions and multi-layer perceptrons to process raw point clouds directly while maintaining permutation invariance. Building upon this foundation, PointNet++ [13] introduces a hierarchical framework that captures local structures through farthest point sampling and multi-scale grouping. Dynamic Graph CNN (DGCNN) [14] further advances local geometric modeling by constructing dynamic graphs on point clouds and applying edge convolution to effectively capture local dependencies. Alongside these foundational approaches, a variety of other architectures have also been developed [15, 16, 17, 39, 19, 20].
@@ 2.2 Backdoor Attacks in 2D Images
While one-to-one attacks are effective, their limitation lies in the inflexible binding between one trigger and one target class. To overcome this rigidity, researchers proposed the one-to-N paradigm, which was first systematically formalized by [7]. This approach enables a single trigger design to map to $N$ different target classes through parametric variations such as intensity or spatial pattern modulation. Beyond this foundational work, the one-to-N concept has been extended to more generalized settings [9, 11, 8, 10], including universal backdoor attacks [8] that target all classes and Marksman backdoors [9] that support arbitrary target selection during inference.
@@ 3.3 Problem Formulation
3.4 One-to-N Attack Framework with Spherical Triggers
Figure 1: Conceptual evolution from static one-to-one to dynamic one-to-N backdoor paradigms in 3D point clouds. (a) Sphere-based one-to-one attack employs a fixed spherical trigger at a static position, functioning as a single key that invariably activates the same target class. (b) Single-sphere design in our one-to-N attack redefines the trigger as a configurable entity, where its spatial position acts as a dynamic parameter to encode different target classes, enabling multiple attacks from a single base trigger. (c) Dual-sphere design in our one-to-N attack further expands the attack capacity by leveraging the combinatorial configuration of two spheres, demonstrating a richer, multi-dimensional parameter space for complex multi-target backdoor manipulations.
Figure 2: End-to-end pipeline of the spherical trigger-based one-to-N attack (single-sphere design). The framework (1) implants class-specific spherical triggers into clean samples; (2) trains a model to associate trigger configurations with target classes while maintaining accuracy on clean inputs; (3) enables multi-target activation during inference (see Fig. 1), where applying a specific trigger causes misclassification to its designated target.
This work establishes the first One-to-N backdoor attack framework for 3D point clouds, introducing a new paradigm, superseding conventional single-target attacks. Our framework systematically coordinates the entire attack lifecycle through seven cohesive phases: (1) Base Entity Selection—choosing a minimal yet effective trigger geometry; (2) Identification of Key Parameters—determining the discriminative features for target control; (3) Trigger Configuration—designing the spatial encoding schemes; (4) Parameter Set Optimization—selecting the optimal $N$ parameter configurations; (5) Poisoned Data Generation—constructing the training corpus with embedded triggers; (6) Model Training—training the model on the poisoned dataset to learn both benign and backdoor mappings; and (7) Multi-Target Activation—executing the attack during inference. The efficacy of this integrated process is supported by both theoretical analysis and empirical validation, as detailed in the subsequent subsections.
@@ 3.4.1 Base Trigger Entity Selection
The one-to-N attack framework associates each target class $t_{i}$ ($1\leq i\leq N$) with a unique trigger configuration $S_{n}$, which is a parametric variation of a base trigger entity. The selection of an appropriate base entity is therefore crucial and should satisfy three key requirements: (1) it must possess a parameter space that can be leveraged to create distinct configurations for different target classes; (2) its design should enhance robustness against pre-processing methods while minimizing perceptibility to human observers; and (3) most importantly, it must be capable of realizing one-to-one attacks, thereby demonstrating the potential for one-to-N attacks through parametric variations.
@@ 3.4.2 Identification of Key Parameters
Experimental Validation. Our experimental investigation provides comprehensive validation of two key theoretical predictions derived from the above analysis: (1) backdoor activation depends critically on the trigger’s spatial position, not merely its presence; and (2) the model’s predictive probability for the target class, i.e., $f_{\bm{\theta}}(\bm{X}^{\prime}_{R})[t_{i}]$, exhibits a decreasing trend with increasing distance from the training position $R_{0}$. Both theoretical propositions are strongly supported by our empirical results, confirming that spatial configuration serves as a powerful and reliable discriminator for target class assignment in our one-to-N attack framework.
@@ 3.4.3 Trigger Configuration
With the base entity and key parameter defined, we now present two concrete trigger implementations—the single-sphere and multi-sphere designs—which are detailed below.
@@ 3.4.4 Parameter Set Optimization
The core objective of parameter set optimization is to select optimal spatial configurations for N triggers within the normalized 3D bounding box, with spatial location serving as the discriminative parameter. To this end, maintaining sufficient separation between triggers is critical for preserving the spatial specificity, as established in our theoretical analysis. This requirement leads to an NP-hard maximin optimization problem aimed at maximizing the minimum pairwise Euclidean distance. We address this challenge by developing efficient greedy algorithms that iteratively select trigger positions from a candidate set to ensure adequate separation. (See Appendix Algorithms 1 and 2 for detailed pseudocode.)
3.4.5 Poisoned Dataset Construction
The poisoned dataset construction process, illustrated in Fig. 2, is formalized as follows. We construct the poisoned dataset by:
@@ 3.4.5 Poisoned Dataset Construction
3.4.6 Model Training
The model is trained following Eq. (4) using poisoned samples from all $N$ targets in a single process, learning the association between spherical configurations and target labels. The resulting model exhibits dual behavior: correct classification of clean inputs and targeted misclassification when triggers are present. This establishes the complete one-to-N capability, enabling selective activation as shown in Fig. 2, where triggers at $(0.95,0.95,0.95)$ and $(0.05,0.05,0.05)$ are mapped to classes Girl and Table, respectively.
3.4.7 Multi-Target Activation
During inference, multi-target activation is accomplished by implanting the pre-configured spherical triggers into clean inputs. The activation mechanism is illustrated in Fig. 1(b) and (c) where, for instance, a single sphere at $(0.95,0.95,0.95)$ leads to a misclassification to girl in a model poisoned with single-sphere triggers, and a dual-sphere pair at $(0.95,0.95,0.95)$ and $(0.95,0.95,0.05)$ results in a misclassification to girl in a model poisoned with dual-sphere triggers.
@@ 4.1.1 Datasets
We evaluate the proposed one-to-N backdoor attack on three widely adopted 3D point cloud benchmarks:
@@ 4.1.2 Models
We evaluate our attack on three representative 3D point cloud classification architectures: PointNet [12], PointNet++ [13], DGCNN [14]. We select the three backbones because of their coverage of fundamental paradigms and established status as benchmarks.
@@ 4.1.3 Evaluation Metrics
We employ two metrics. The first one is Attack Success Rate (ASR), which is defined as the percentage of triggered samples from non-target classes that are misclassified into the target labels. ASR evaluates the effectiveness of the backdoor attack. The second metric is Accuracy (ACC), which is the standard classification accuracy of the model on a clean test set. It evaluates the model’s performance on clean inputs, ensuring that the backdoor attack does not compromise its normal functionality. ACC serves as an important indicator of attack stealthiness: the smaller the degradation in ACC after dataset poisoning, the more covert and effective the backdoor attack. Beyond ACC measurements, stealthiness can be gauged through the metric of perceptual resistance, i.e., a trigger’s ability to evade visual detection and pre-processing defense mechanisms.
@@ 4.2.1 Empirical Evidence of Spatial Specificity
This overall decreasing trend provides the foundational principle for trigger placement in one-to-N attacks. The existence of such a relationship implies that, for any chosen ASR threshold defining acceptable interference, there exists a corresponding critical distance. In our specific experimental configuration, an ASR threshold of 5% corresponds to an observed critical distance of approximately 0.26. This empirical observation demonstrates that to prevent significant interference between any two triggers in the one-to-N framework, their separation must exceed the critical distance specific to the model and data distribution. Therefore, our objective of maximizing the minimum inter-trigger distance is directly motivated by the need to mitigate interference between triggers, thereby ensuring reliable discrimination among all target classes in the one-to-N attack.
@@ 4.2.2 Impact of Poisoning Ratio
This experiment evaluates the impact of poisoning ratio on our one-to-N backdoor attack under a fixed configuration of 4 target classes using the PointNet model. The primary objective is to analyze the performance trends across varying poisoning ratios (0.3%–4%) to identify an effective range for subsequent experiments, while providing initial observations on the comparative efficacy of single-sphere versus dual-sphere triggers and their performance across different datasets.
@@
Notably, the performance gains diminish significantly beyond the 2% poisoning threshold, suggesting an operational sweet spot for practical attack deployment. At a minimal overall poisoning ratio of 0.3% (equivalent to only 0.075% per target for N=4), the attack already achieves substantial success rates, reaching up to 90% on some datasets and maintaining above 50% even in the most challenging configurations. This highlights the efficiency of our approach in establishing multiple backdoors under severely constrained poisoning budgets.

### v2

# STONE: Pioneering the One-to-N Backdoor Threat in 3D Point Cloud

@@ Abstract
Backdoor attacks pose a critical threat to deep learning, especially in safety-sensitive 3D domains such as autonomous driving and robotics. Despite their potency, existing attacks on 3D point clouds are limited to a static one-to-one paradigm, leaving the more flexible one-to-N backdoor threat largely unexplored and without a theoretical or practical foundation. We address this by introducing STONE (Spherical Trigger One-to-N Backdoor Enabling), the first framework that instantiates this threat through a configurable spherical trigger. Its parameterizable spatial properties create a dynamic key space, enabling a single trigger to control multiple output labels. Theoretically, we ground STONE through Neural Tangent Kernel (NTK) analysis, providing the first formal basis for one-to-N mappings in 3D models. Empirically, extensive evaluations show high attack success rate (up to 100%) with no loss in clean-data accuracy. This work establishes a foundational benchmark for multi-target threats in 3D vision, crucial for securing future intelligent systems.
@@ 1 Introduction
Current research on 3D backdoor attacks has primarily focused on one-to-one attack paradigms, where a single trigger pattern is designed to cause the model to misclassify any input containing it as a specific target class determined during the training phase. Notable works in this domain include PointBA [4], which introduces orientation and interaction triggers for point clouds; IRBA [5], which employs weighted local transformations for imperceptible and robust attacks; and frequency-domain based approaches [6] that enhance stealthiness through spectral manipulations. While these methods have demonstrated impressive attack success rates, they share a fundamental limitation: the attack is confined at inference to misclassifying inputs into the sole target class determined during training. This constraint significantly reduces the attack flexibility and practicality in real-world scenarios where attackers may need to target multiple classes simultaneously.
@@
The concept of one-to-N backdoor attacks was first systematically formalized in the work by Xue et al. [7] for 2D image domains, representing a significant evolution beyond traditional one-to-one paradigms. This approach enables a single trigger design to map to $N$ target classes through parametric variations such as its intensity or spatial pattern. This contrasts sharply with the N-to-N paradigm of combining multiple distinct trigger types, which suffers from scalability limitations stemming from the finite number of available types, and performance inconsistencies due to divergent optimization requirements. Building upon this foundational concept, the one-to-N principle has been extended to more generalized settings [8, 9, 10, 11, 12], including universal backdoor attacks [8] targeting all classes and Marksman backdoors [9] supporting arbitrary target selection during inference. However, the translation of this powerful one-to-N paradigm into the 3D domain presents significant challenges that remain. The structural transition from dense, grid-aligned pixels to unstructured point clouds invalidates core assumptions of 2D trigger design and demands a novel instantiation of the one-to-N principle. Consequently, while 3D backdoor research has advanced in one-to-one mappings, the potential for one-to-N attacks has remained largely unexplored due to these inherent structural barriers and the absence of both theoretical foundation and practical implementation.
To address this critical gap, we introduce STONE (Spherical Trigger One-to-N Backdoor Enabling), the first comprehensive framework that establishes and instantiates the one-to-N backdoor threat in 3D point clouds. Our approach leverages a configurable spherical trigger whose spatial properties form a dynamic key space, enabling a single trigger instance to map to diverse target labels. We ground this capability theoretically through Neural Tangent Kernel (NTK) analysis, providing the first formal basis for one-to-N mappings in 3D models, with extensive experimental validation across multiple datasets and architectures confirming both high attack success and preserved clean-data performance.
The main contributions of this work are:
@@
    The first one-to-N backdoor framework for 3D point clouds: We establish and instantiate this advanced threat paradigm, moving beyond the limitations of traditional one-to-one attacks.
@@
The remainder of this paper is structured as follows. Section 2 reviews related work on 3D point cloud deep learning, backdoor attacks in both 2D and 3D domains, and existing defense mechanisms. Section 3 presents our proposed STONE framework, detailing the problem formulation, spherical trigger design with parameterizable spatial properties, and the complete attack pipeline. Section 4 provides extensive experimental validation across multiple datasets and model architectures, evaluating attack effectiveness, scalability with varying numbers of target classes, and resistance to defenses. Section 5 summarizes our key contributions and discusses future research directions. The Appendix contains supporting theoretical analysis using Neural Tangent Kernel theory and detailed trigger placement algorithms.
@@ 2.1 Deep Learning on 3D Point Cloud
The evolution of direct point cloud processing networks has addressed these challenges through several representative architectures. PointNet [13] pioneers the use of symmetric functions and multi-layer perceptrons to process raw point clouds directly while maintaining permutation invariance. Building upon this foundation, PointNet++ [14] introduces a hierarchical framework that captures local structures through farthest point sampling and multi-scale grouping. Dynamic Graph CNN (DGCNN) [15] further advances local geometric modeling by constructing dynamic graphs on point clouds and applying edge convolution to effectively capture local dependencies. Alongside these foundational approaches, a variety of other architectures have also been developed [16, 17, 18, 19, 20].
@@ 2.2 Backdoor Attacks in 2D Images
While one-to-one attacks are effective, their limitation lies in the inflexible binding between one trigger and one target class. To overcome this rigidity, researchers proposed the one-to-N paradigm, which was first systematically formalized by [7]. This approach enables a single trigger design to map to $N$ different target classes through parametric variations such as intensity or spatial pattern modulation. Beyond this foundational work, the one-to-N concept has been extended to more generalized settings [9, 11, 8, 10, 12], including universal backdoor attacks [8] that target all classes and Marksman backdoors [9] that support arbitrary target selection during inference.
@@ 3.3 Problem Formulation
3.4 STONE: One-to-N Backdoor Framework with Spherical Triggers
Figure 1: Conceptual evolution from static one-to-one to dynamic one-to-N backdoor paradigms in 3D point clouds. (a) Traditional one-to-one attack employs a fixed spherical trigger, functioning as a single key that activates the same target class. (b) Our STONE framework redefines the trigger as a configurable entity in the single-sphere design, where spatial position acts as a dynamic parameter to encode different target classes. (c) The dual-sphere design further expands capacity through combinatorial configurations, demonstrating a richer parameter space for complex multi-target manipulations.
Figure 2: End-to-end pipeline of the STONE framework (single-sphere design). The framework (1) implants class-specific spherical triggers into clean samples; (2) trains a model to associate trigger configurations with target classes while maintaining accuracy on clean inputs; (3) enables multi-target activation during inference, where applying a specific trigger causes misclassification to its designated target.
We present STONE (Spherical Trigger One-to-N Backdoor Enabling), the first comprehensive framework that establishes and instantiates the one-to-N backdoor threat in 3D point clouds. Our framework systematically coordinates the attack lifecycle through seven cohesive phases: (1) Base Entity Selection—choosing an effective trigger geometry; (2) Identification of Key Parameters—determining discriminative features for target control; (3) Trigger Configuration—designing spatial encoding schemes; (4) Parameter Set Optimization—selecting optimal parameter configurations; (5) Poisoned Data Generation—constructing the training corpus with embedded triggers; (6) Model Training—learning both benign and backdoor mappings; and (7) Multi-Target Activation—executing dynamic attacks during inference.
@@ 3.4.1 Base Trigger Entity Selection
The core insight of STONE is to associate each target class $t_{i}$ ($1\leq i\leq N$) with a unique configuration $S_{n}$, derived through parametric variations of a base trigger entity. The selection of this base entity is crucial and must satisfy: (1) possessing a parameterizable space for creating distinct configurations; (2) maintaining stealthiness against detection; and (3) demonstrating proven effectiveness in basic one-to-one attacks to ensure learnability.
@@ 3.4.2 Identification of Key Parameters
Experimental Validation. We conducted experiments to validate two key theoretical predictions: (1) backdoor activation is highly sensitive to the trigger’s spatial position, not just its presence; and (2) the model’s predictive probability for the target class, $f_{\bm{\theta}}(\bm{X}^{\prime}_{R})[t_{i}]$, exhibits a negative correlation with the distance from the training position $R_{0}$. The empirical results confirm these hypotheses, demonstrating that spatial configuration is a powerful and reliable feature for discriminating target classes in our one-to-N attack.
@@ 3.4.3 Trigger Configuration
We implement two concrete configurations within the STONE framework:
@@ 3.4.4 Parameter Set Optimization
The core objective of parameter set optimization is to select optimal configurations for $N$ triggers within the normalized 3D bounding box, with spatial location serving as the discriminative parameter. To this end, maintaining sufficient separation between triggers is critical for preserving the spatial specificity, as established in our theoretical analysis. This requirement leads to an NP-hard maximin optimization problem aimed at maximizing the minimum pairwise Euclidean distance. We address this challenge by developing efficient greedy algorithms that iteratively select trigger positions from a candidate set to ensure adequate separation. (See Appendix Algorithms 1 and 2 for detailed pseudocode.)
3.4.5 End-to-End Attack Pipeline
The STONE framework completes the attack lifecycle through three final phases: Poisoned Dataset Construction, Model Training, and Multi-Target Activation, which together form an end-to-end pipeline for executing the one-to-N backdoor attack.
Poisoned Dataset Construction. The poisoned dataset construction process, illustrated in Fig. 2, is formalized as follows. We construct the poisoned dataset by:
@@ 3.4.5 End-to-End Attack Pipeline
Model Training. The model is trained following Eq. (4) using poisoned samples from all $N$ targets in a single process, learning the association between spherical configurations and target labels. The resulting model exhibits dual behavior: correct classification of clean inputs and targeted misclassification when triggers are present. This establishes the complete one-to-N capability, enabling selective activation as shown in Fig. 2, where triggers at $(0.95,0.95,0.95)$ and $(0.05,0.05,0.05)$ are mapped to classes Girl and Table, respectively.
Multi-Target Activation. During inference, multi-target activation is accomplished by implanting the pre-configured spherical triggers into clean inputs. The activation mechanism is illustrated in Fig. 1(b) and (c) where, for instance, a single sphere at $(0.95,0.95,0.95)$ leads to a misclassification to girl in a model poisoned with single-sphere triggers, and a dual-sphere pair at $(0.95,0.95,0.95)$ and $(0.95,0.95,0.05)$ results in a misclassification to girl in a model poisoned with dual-sphere triggers.
@@ 4.1.1 Datasets
We conduct extensive evaluations of the proposed STONE framework across multiple widely adopted 3D point cloud benchmarks to ensure comprehensive validation:
@@ 4.1.2 Models
To validate the generalizability of our approach, we evaluate STONE on three representative 3D deep learning architectures: PointNet [13], PointNet++ [14], and DGCNN [15]. These models cover fundamental paradigms including point-wise MLPs, hierarchical feature learning, and graph-based convolutions, providing a robust testbed for evaluating the proposed one-to-N backdoor attacks across different learning mechanisms.
@@ 4.1.3 Evaluation Metrics
We employ two primary metrics to comprehensively assess STONE’s performance. The first one is Attack Success Rate (ASR), which is defined as the percentage of triggered samples from non-target classes that are misclassified into the target labels. ASR evaluates the effectiveness of the backdoor attack. The second metric is Accuracy (ACC), which is the standard classification accuracy of the model on a clean test set. It evaluates the model’s performance on clean inputs, ensuring that the backdoor attack does not compromise its normal functionality. ACC serves as an important indicator of attack stealthiness: the smaller the degradation in ACC after dataset poisoning, the more covert and effective the backdoor attack. Beyond ACC measurements, stealthiness can be gauged through the metric of perceptual resistance, i.e., a trigger’s ability to evade visual detection and pre-processing defense mechanisms.
@@ 4.2.1 Empirical Evidence of Spatial Specificity
This overall decreasing trend provides the foundational principle for trigger placement in the STONE framework. The existence of such a relationship implies that, for any chosen ASR threshold defining acceptable interference, there exists a corresponding critical distance. In our specific experimental configuration, an ASR threshold of 5% corresponds to an observed critical distance of approximately 0.26. This empirical observation demonstrates that to prevent significant interference between any two triggers in the one-to-N framework, their separation must exceed the critical distance specific to the model and data distribution. Therefore, our objective of maximizing the minimum inter-trigger distance is directly motivated by the need to mitigate interference between triggers, thereby ensuring reliable discrimination among all target classes in the STONE framework..
@@ 4.2.2 Impact of Poisoning Ratio
This experiment evaluates the impact of poisoning ratio on the STONE framework under a fixed configuration of 4 target classes using the PointNet model. The primary objective is to analyze the performance trends across varying poisoning ratios (0.3%–4%) to identify an effective range for subsequent experiments, while providing initial observations on the comparative efficacy of single-sphere versus dual-sphere triggers and their performance across different datasets.
@@
Notably, the performance gains diminish significantly beyond the 2% poisoning threshold, suggesting an operational sweet spot for practical attack deployment. At a minimal overall poisoning ratio of 0.3% (equivalent to only 0.075% per target for N=4), the attack already achieves substantial success rates, reaching up to 90% on some datasets and maintaining above 50% even in the most challenging configurations. This highlights the efficiency of the STONE framework in establishing multiple backdoors under severely constrained poisoning budgets.